# MIXRTs: Toward Interpretable Multi-Agent Reinforcement Learning via Mixing Recurrent Soft Decision Trees

### v3

@@ Abstract
While achieving tremendous success in various fields, existing multi-agent reinforcement learning (MARL) with a black-box neural network architecture makes decisions in an opaque manner that hinders humans from understanding the learned knowledge and how input observations influence decisions. Instead, existing interpretable approaches, such as traditional linear models and decision trees, usually suffer from weak expressivity and low accuracy. To address this apparent dichotomy between performance and interpretability, our solution, MIXing Recurrent soft decision Trees (MIXRTs), is a novel interpretable architecture that can represent explicit decision processes via the root-to-leaf path and reflect each agent’s contribution to the team. Specifically, we construct a novel soft decision tree to address partial observability by leveraging the advances in recurrent neural networks, and demonstrate which features influence the decision-making process through the tree-based model. Then, based on the value decomposition framework, we linearly assign credit to each agent by explicitly mixing individual action values to estimate the joint action value using only local observations, providing new insights into how agents cooperate to accomplish the task. Theoretical analysis shows that MIXRTs guarantees the structural constraint on additivity and monotonicity in the factorization of joint action values. Evaluations on the challenging Spread and StarCraft II tasks show that MIXRTs achieves competitive performance compared to widely investigated methods and delivers more straightforward explanations of the decision processes. We explore a promising path toward developing learning algorithms with both high performance and interpretability, potentially shedding light on new interpretable paradigms for MARL.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has proven to be a considerable promise for addressing a variety of challenging tasks, e.g., games [1, 2, 3], autonomous driving [4, 5] and robotics interactions [6, 7]. Despite its promising performance, the advancements are mainly attributed to utilizing deep neural network (DNN) models as powerful function approximators, which are encoded with thousands to millions of parameters interacting in complex and nonlinear ways [8]. This complexity of architecture brings substantial obstacles for humans to understand how decisions are made and what key features influence decisions, especially when the network grows deeper in size or more complex structures are appended [9, 10]. Indeed, creating mechanisms to interpret the implicit behaviors of black-box DNNs remains an open problem in the field of machine learning [11, 12].
@@
It is crucial to gain insights into the decision-making process of artificially intelligent agents for successful and reliable deployments into real-world applications, especially in high-stakes domains such as healthcare, aviation, and public policy making [13, 14, 15, 16]. Significant limitations are imposed on practitioners applying MARL techniques, as the lack of transparency creates critical barriers to establishing trust in the learned policies and scrutinizing policy weaknesses [17]. Explainable reinforcement learning [18, 19, 20] stands as a promising avenue for the formulation of transparent procedures that can be followed step-by-step or held accountable by human operators. While existing explainable methods offer some potential [21, 22] or provide vision-based interpretations [23] in single-agent tasks, they are still struggling to strike a balance between trustworthy model interpretability and matching the performance in look-ahead reinforcement learning [24], especially in multi-agent domains.
@@
Traditional decision trees [25, 26] present lightweight inferences to interpretability, as humans can easily understand the decision process by visualizing decision pathways. However, traditional decision trees may suffer from weak expressivity and therefore low accuracy with a shallow tree depth and univariate decision nodes, making it a hard trade-off between model interpretability and learning performance. Alternatively, differentiable soft decision trees (SDTs) [27, 28], based on the structure of fuzzy decision trees, model expressivity lies in the middle of traditional decision trees and neural networks [29]. SDTs are acclaimed for their comprehensibility, offering interpretations that non-expert individuals can easily visualize and simulate, thereby enhancing human readability. Several works [30, 31] attempt to utilize an imitation learning paradigm to distill a pre-trained DNN control policy into an SDT, providing an interpretable form of policy in single-agent tasks. However, the inherent opacity of DNN structures imposes constraints on the SDT’s capacity to accurately mimic the DNN policy. It thus poses a challenge to jointly maintaining the mimic fidelity and the simplicity of tree models to find a “sweet spot”.
@@
Rather than imitating a pre-trained policy, an alternative is to train an SDT policy from agent experience in an end-to-end manner, which directly learns the domain knowledge from the tasks using interpretable models. While SDT has achieved an adequate balance between interpretability and performance in simple single-agent domains (e.g., CartPole and MountainCar in OpenAI Gym [32][31], they cannot maintain satisfying learning performance without sacrificing interpretability in complex multi-agent tasks due to limited model expressivity. Particularly, to the best of our knowledge, there is seldom existing work exploring the model interpretability in MARL domains. Here, we aim at a more efficient and explicit interpretable architecture with tree-based models in multi-agent challenging tasks.
@@
In this paper, we propose a novel MIXing Recurrent soft decision Trees (MIXRTs) method to tackle the tension between model interpretability and learning performance in MARL domains. Instead of trying to understand how a DNN makes its decisions, we utilize the differentiable SDT to keep the decision-making process. First, to facilitate learning over long timescales for each agent, we propose the recurrent tree cell (RTC) that receives the current individual observation and history information as input at each timestamp. Second, we utilize a linear combination of multiple RTCs with an ensemble mechanism to improve performance and reduce the variance of SDTs, while maintaining the interpretability of RTCs. By visualizing the tree structure, RTCs can provide intuitive explanations of how important features affect the decision process. MIXRTs consists of individual RTCs representing the individual value function, and a mixing tree structure aiming to learn an optimal linear value decomposition, which ensures consistency between the centralized and decentralized policies. The linear mixing structure emphasizes the explanations about what role each agent plays in cooperative tasks by analyzing its assigned credit. To improve learning efficiency, we utilize parameters sharing across individual RTCs to dramatically reduce the number of policy parameters, while experience can be shared across other agents. We evaluate MIXRTs on a range of challenging tasks in Spread and StarCraft II [33] environment. Empirical results substantiate that our learning architecture delivers simple explanations while enjoying competitive performance. Specifically, MIXRTs finds desirable optimal policies in easy scenarios compared to the baselines like the widely investigated QMIX [3] and QPLEX [34].
@@
The remaining paper is organized as follows. In Section 2, we introduce basic concepts of MARL, SDTs, and related work. In Section 3 and 4, we present the RTCs and the linear mixing architecture of individual RTCs, respectively. In Section 5, we give experimental results of learning performance compared to the existent baselines. The comprehensive interpretability of our model and the results of the user study are given in Section 6. Finally, we give concluding remarks in Section 7.
@@ 2.1 Preliminaries
Dec-POMDP. The fully cooperative multi-agent task is generally modeled as a decentralized partially observable Markov decision process (Dec-POMDP) [35] that consists of a tuple $\mathcal{S},\mathcal{U},\mathcal{P},r,\mathcal{Z},\mathcal{O},n,\gamma\right>$, where $s\in\mathcal{S}$ describes the global state of the environment. At each time step, each agent $i\in\{1,...,n\}$ only receives a partial observation $o_{i}\in\mathcal{Z}$ from an observation function $\mathcal{O}(s,i)$, and chooses an action $u_{i}\in\mathcal{U}$ to formulate a joint action $\bm{u}:=\left[u_{i}\right]_{i=1}^{n}\in\mathcal{U}^{n}$. This results in a transition to next state $s^{\prime}\sim P\left(s^{\prime}\mid s,\bm{u}\right)$. All agents share the same team reward signal $r(s,\bm{u}):\mathcal{S}\times\mathcal{U}^{n}\rightarrow\mathbb{R}$. Furthermore, each agent learns its own policy $\pi_{i}\left(u_{i}\mid\tau_{i}\right):\mathcal{T}\times\mathcal{U}\rightarrow[0,1]$, which conditions on the its action-observation history denoted as $\tau_{i}\in\mathcal{T}:=(\mathcal{Z}\times\mathcal{U})^{*}$. The goal is to find an optimal joint policy $\bm{\pi}=\left\langle\pi_{1},\ldots,\pi_{n}\right\rangle$ to maximize the discounted cumulative return $\sum_{t=0}^{\infty}\gamma^{t}r^{t}$, where the discount factor is $\gamma\in[0,1)$.
@@
Multi-agent Q-learning in Dec-POMDP. Q-learning [36] is a classic tabular model-free algorithm to find the optimal joint action-value function $Q^{*}(s,\bm{u})=r(s,\bm{u})+\gamma\mathbb{E}_{s^{\prime}}[\max_{\bm{u}^{\prime}\in\mathcal{U}^{n}}$ $Q^{*}\left(s^{\prime},\bm{u}^{\prime}\right)]$. Multi-agent Q-learning approaches [2, 37, 38] are based on the value decomposition extension of deep Q-learning [36, 39], thus the agent system receives the joint action-observation history $\bm{\tau}$ and the joint action $\bm{u}$. Given transition tuples $(\bm{\tau},\bm{u},r,\bm{\tau^{\prime}})$ from the replay buffer $\mathcal{B}$, the network parameters $\mathit{\Theta}$ are learnt by minimizing the squared loss $\mathcal{L}(\mathit{\Theta})$ on the temporal-difference (TD) error $\sigma=y^{\prime}-Q_{tot}(\bm{\tau},\bm{u};\mathit{\Theta})$, where $y^{\prime}=r+\gamma\,\max_{\bm{u^{\prime}}}Q_{tot}(\bm{\tau^{\prime}},\bm{u^{\prime}};\mathit{\Theta^{\prime}})$ is the target and $Q_{tot}(\bm{\tau},\bm{u};\mathit{\Theta})$ is used in place of $Q_{tot}(s,\bm{u};\mathit{\Theta})$ due to partial observability. The parameters $\mathit{\Theta^{\prime}}$ from a target network are periodically copied from $\mathit{\Theta}$ and remain constant over multiple iterations.
@@
Centralized training and decentralized execution (CTDE). In the CTDE fashion, action-observation values of all agents and global states are collected through a central controller during training. However, in the execution phase, each agent has its own policy network to make decisions based on its individual action-observation history. The individual-global-max (IGM) property [37] is a popular principle to realize effective CTDE as
@@
where $\bm{\tau}$ represents the history record of joint action-observation and $Q_{tot}$ represents the joint Q-function for each individual action-value function $Q_{i}(\tau_{i},u_{i})$. Both VDN [2] and QMIX [3] are popular CTDE methods for estimating the optimal $Q_{tot}$ via the combined agents of $Q_{i}$. In VDN, the $Q_{tot}$ is calculated by summing the utilities of each agent as $Q_{tot}(\bm{\tau},\bm{u})=\sum_{i=1}^{n}Q_{i}(\tau_{i},u_{i})$. QMIX combines $Q_{i}$ through the state-dependent nonlinear monotonic function $f_{s}$: $Q_{tot}(\bm{\tau},\bm{u})=f_{s}(Q_{1}(\tau_{1},u_{1}),...,Q_{i}(\tau_{i},u_{i}),...,Q_{n}(\tau_{n},u_{n}))$, where $\frac{\partial f_{s}}{\partial Q_{i}}\geq 0,\forall i\in\{1,...,n\}$.
Soft decision trees (SDTs). Differentiable decision trees [40, 41] have been shown to be more accurate compared to traditional hard decision trees. Especially, an SDT [27] with binary probabilistic decision boundary at each node has favorable transparency and performance in reinforcement learning tasks [30, 31]. As shown in Fig. 1(a), SDT fixes the depth of the tree in advance and performs path selection by probability, which is different from the traditional growth way of learning crisp decision boundaries. By giving an observation $o^{t}$, each inner node $j$ calculates the probability for traversing to the left child node as
@@
Figure 1: Examples of SDT and RTC. (a) A two-level SDT. (b) Illustration of the process of a two-level RTC that receives the current individual observation $o^{t}_{i}$ and the historical leaf information $h^{t-1}_{i}$ as input at each timestamp.
@@ 2.2 Related works
Explainable reinforcement learning. In the reinforcement learning community, a number of prior works have started exploring approaches that lead to interpretable policies via intrinsic methods [28, 40, 41, 42]. Intrinsic methods generally rely on using inherently interpretable models such as classical decision trees and linear models, which directly represent the faithful decision-making process. Conservative Q-improvement method [42] presents an interpretable decision-tree model that learns a policy in the form of a hard decision tree for the robot navigation tasks. It can better trade off the performance and simplicity of the policy model by adding a new split only if doing so could improve the policy. Differentiable decision trees methods [28, 43] allow for a gradient update over Q-learning and policy gradient algorithms by replacing the Boolean decision with the sigmoid activation function at the decision node, which improves the performance while affording interpretable tree-based policies. Rather than incrementally adding nodes, differentiable decision trees produce the trees entirely over the learning process, where the leaf nodes represent single features with the discretization technique for improving interpretability. Instead of constructing a univariate decision tree, cascading decision tree (CDT) [31], as an extension based on SDT with multivariate decision nodes, applies representation learning to the decision path to allow richer expressivity. The prior works closest to ours are SDT [27] and CDT [31], which policies are generated by the routing probabilities at each leaf node to make decisions. Our method differs in that 1) we introduce the recurrency into SDTs for capturing the long-term condition in partially observable tasks, 2) we explicitly present the graphical decision process with less depth by considering all variate at each node instead of univariate feature, and 3) we interpret the role each agent plays in the team by visualizing the linear credit assignment.
@@
Cooperative multi-agent reinforcement learning. Dating back to the early stage of MARL, independent Q-learning [44] is the most commonly used method where each agent myopically optimizes its individual policy that is independent of the others. Generally, they only observe local information, execute their own actions, and receive their rewards individually such that other agents are perceived as part of the environment. This approach has no theoretical convergence guarantee like Q-learning for ignoring the nonstationarity induced by the changes of the other agents. By contrast, CTDE [45, 46] is an advanced paradigm for cooperative MARL tasks, which allows each agent to learn their individual utility functions via optimizing the joint action-value function for credit assignment. VDN [2] factorizes the joint action-value function into a linear summation over individual agents. To overcome the VDN’s limitations that ignore extra global state information, QMIX [3] utilizes a network to estimate joint action-value function with a non-linear combination of each agent value. QTRAN [37] relaxes these constraints of the greedy action selections between the joint and individual value functions with two soft regularizations. Different from QTRAN which loses the exact IGM consistency, QPLEX [34] uses a dueling network architecture to represent both joint and individual action-value functions to guarantee the IGM property.
Our proposed MIXRTs is distinguished from the above representative value decomposition methods in the following ways: 1) our method builds upon soft decision trees; our RTCs not only produces an intrinsically interpretable model for visualizing the decision-making process instead of the complex neural network architecture but also achieve competitive performance. 2) Our mixing trees module linearly factorizes the joint action-value function into individual action-value, achieving high scalability with a full realization of the IGM principle. Benefiting from the interpretable model, we can explain not only the explicit behavior of the individual agent but also the behavior of different roles in the allies. 3) Our artful designs yield lightweight inference, whilst achieving competitive performance on a range of cooperative challenging tasks.

### v4

@@ Abstract
While achieving tremendous success in various fields, existing multi-agent reinforcement learning (MARL) with a black-box neural network makes decisions in an opaque manner that hinders humans from understanding the learned knowledge and how input observations influence decisions. In contrast, existing interpretable approaches usually suffer from weak expressivity and low performance. To bridge this gap, we propose MIXing Recurrent soft decision Trees (MIXRTs), a novel interpretable architecture that can represent explicit decision processes via the root-to-leaf path and reflect each agent’s contribution to the team. Specifically, we construct a novel soft decision tree using a recurrent structure and demonstrate which features influence the decision-making process. Then, based on the value decomposition framework, we linearly assign credit to each agent by explicitly mixing individual action values to estimate the joint action value using only local observations, providing new insights into interpreting the cooperation mechanism. Theoretical analysis confirms that MIXRTs guarantee additivity and monotonicity in the factorization of joint action values. Evaluations on complex tasks like Spread and StarCraft II demonstrate that MIXRTs compete with existing methods while providing clear explanations, paving the way for interpretable and high-performing MARL systems.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has been shown considerable potential in solving a variety of challenging tasks, e.g., games [1, 2, 3], autonomous driving [4, 5, 6] and robotics interactions [7, 8]. Despite these promising results, much of this progress relies on deep neural network (DNN) models serving as powerful function approximators, encoded with thousands to millions of parameters interacting in complex and nonlinear ways [9]. This architectural complexity brings substantial obstacles for humans to understand how decisions are made and what key features influence decisions, especially when the network becomes deeper in size or more complex structures are appended [10, 11]. Indeed, creating mechanisms to interpret the implicit behaviors of black-box DNNs remains an open problem in the field of machine learning [12, 13].
@@
It is crucial to gain insights into the decision-making process of artificially intelligent agents for their successful and reliable deployment into real-world applications, especially in high-risk domains such as healthcare and military [14, 15, 16, 17]. The lack of transparency in MARL techniques imposes significant limitations on practitioners, bringing critical barriers to establishing trust in the learned policies and scrutinizing their weaknesses [18]. Explainable reinforcement learning [19, 20, 21] emerges as a promising approach to develop transparent procedures that can be followed step-by-step or held accountable by human operators. While existing explainable methods offer some potential [22, 23] or vision-based interpretations [24, 21] in single-agent tasks, they still struggle to balance interpretability and performance in complex reinforcement learning tasks [25], especially in multi-agent domains.
@@
Traditional decision trees [26, 27] provide interpretable inferences at a rule level, as humans can easily understand their decision process by visualizing decision paths. However, they often suffer from limited expressivity and low accuracy when using shallow trees and univariate decision nodes, leading to a difficult trade-off between model interpretability and performance. Alternatively, differentiable soft decision trees (SDTs) [28, 29] are built on the structure of fuzzy decision trees, bridging traditional decision trees and neural networks in terms of expressiveness [30]. SDTs offer interpretations that non-expert individuals can easily visualize and simulate, thereby enhancing human readability. Several works [31, 32] have attempted to use an imitation learning paradigm to distill a pre-trained DNN control policy into an SDT, providing an interpretable form of policy in single-agent tasks. However, the simple structure of an SDT often makes it difficult to accurately mimic its original policy. Therefore, it remains a challenge to find a “sweet spot” that can well balance fidelity and simplicity in tree models.
@@
Rather than imitating a pre-trained policy, an alternative paradigm is to train an SDT policy from agent experience in an end-to-end manner, which directly learns the domain knowledge from tasks using interpretable models. While SDT has achieved an adequate balance between interpretability and performance in simple single-agent domains (e.g., CartPole and MountainCar in OpenAI Gym [33][32], they cannot maintain satisfying learning performance without sacrificing interpretability in complex multi-agent tasks due to limited model expressivity. Particularly, to the best of our knowledge, there is seldom existing work exploring the model interpretability in MARL domains. Here, our goal is to strike a favorable balance between interpretability and insight into the underlying decision-making process using tree-based models in a multi-agent system.
@@
In this paper, we propose a novel MIXing Recurrent soft decision Trees (MIXRTs) method to tackle the tension between model interpretability and learning performance in MARL domains. Instead of attempting to understand how a DNN makes its decisions, we utilize a differentiable SDT to learn the decision-making process for a certain task. First, to facilitate learning over long timescales for each agent, we propose the recurrent tree cell (RTC) that receives the current individual observation and history information as input at each timestamp. Second, we utilize a linear combination of multiple RTCs to improve performance and reduce the variance of SDTs, while maintaining the interpretability of RTCs. By visualizing the tree structure, RTCs can provide intuitive explanations of how important features affect the decision process. MIXRTs consist of RTCs representing the individual value function and a mixing tree structure aiming to learn an optimal linear value decomposition, which ensures consistency between the centralized and decentralized policies. The linear mixing structure emphasizes the explanations about what role each agent plays in cooperative tasks by analyzing its assigned credit. To improve learning efficiency, we also use parameter sharing across individual RTCs to dramatically reduce the number of policy parameters, therefore experience can be shared across other agents. We evaluate MIXRTs on a range of challenging tasks in Spread and StarCraft II [34] environment. Empirical results show that our learning architecture delivers simple explanations while enjoying competitive performance. Specifically, MIXRTs find desirable optimal policies in easy scenarios compared to popular baselines like the widely investigated QMIX [3] and QPLEX [35].
@@
The remaining paper is organized as follows. In Section 2, we introduce basic concepts of MARL, SDTs, and related work. In Section 3 and 4, we present the RTCs and the linear mixing architecture of individual RTCs, respectively. In Section 5, we give experimental results of learning performance compared to the existent baselines. The comprehensive interpretability of our model and results of the user study are given in Section 6. Finally, we give concluding remarks in Section 7.
@@ 2.1 Preliminaries
Dec-POMDP. The fully cooperative multi-agent task is generally modeled as a decentralized partially observable Markov decision process (Dec-POMDP) [36] that consists of a tuple $\mathcal{S},\mathcal{U},\mathcal{P},r,\mathcal{Z},\mathcal{O},n,\gamma\right>$, where $s\in\mathcal{S}$ describes the global state of the environment. At each time step, each agent $i\in\{1,...,n\}$ only receives a partial observation $z\in\mathcal{Z}$ generated from an observation function $\mathcal{O}(s,i):S\times n\rightarrow\mathcal{Z}$, and chooses an action $u_{i}\in\mathcal{U}$ to formulate a joint action $\bm{u}:=\left[u_{i}\right]_{i=1}^{n}\in\mathcal{U}^{n}$. This results in a transition to next state $s^{\prime}\sim P\left(s^{\prime}\mid s,\bm{u}\right)$. All agents share the same team reward signal $r(s,\bm{u}):\mathcal{S}\times\mathcal{U}^{n}\rightarrow\mathbb{R}$. Furthermore, each agent learns its own policy $\pi_{i}\left(u_{i}\mid\tau_{i}\right):\mathcal{T}\times\mathcal{U}\rightarrow[0,1]$, which conditions on its action-observation history $\tau_{i}\in\mathcal{T}$, and we define that $\bm{\tau}\in\bm{\mathcal{T}}$ is the joint action-observation history. The goal is to find an optimal joint policy $\bm{\pi}=\left\langle\pi_{1},\ldots,\pi_{n}\right\rangle$ to maximize the discounted cumulative return $\sum_{t=0}^{\infty}\gamma^{t}r^{t}$, where $\gamma\in[0,1)$ is a discount factor.
@@
Multi-agent Q-learning in Dec-POMDP. Q-learning [37] is a classic tabular model-free algorithm to find the optimal joint action-value function $Q^{*}(s,\bm{u})=r(s,\bm{u})+\gamma\mathbb{E}_{s^{\prime}}[\max_{\bm{u}^{\prime}\in\mathcal{U}^{n}}$ $Q^{*}\left(s^{\prime},\bm{u}^{\prime}\right)]$. Multi-agent Q-learning approaches [2, 38, 39] are generally based on the value decomposition extension of deep Q-learning [37, 40], where the agent system receives the joint action-observation history $\bm{\tau}$ and the joint action $\bm{u}$. Given transition tuples $(\bm{\tau},\bm{u},r,\bm{\tau^{\prime}})$ from the replay buffer $\mathcal{B}$, the network parameters $\mathit{\Theta}$ are learnt by minimizing the squared loss $\mathcal{L}(\mathit{\Theta})$ on the temporal-difference (TD) error $\sigma=y^{\prime}-Q_{tot}(\bm{\tau},\bm{u};\mathit{\Theta})$, where $y^{\prime}=r+\gamma\,\max_{\bm{u^{\prime}}}Q_{tot}(\bm{\tau^{\prime}},\bm{u^{\prime}};\mathit{\Theta^{\prime}})$ is the target and $Q_{tot}(\bm{\tau},\bm{u};\mathit{\Theta})$ is used in place of $Q_{tot}(s,\bm{u};\mathit{\Theta})$ due to partial observability. The parameters $\mathit{\Theta^{\prime}}$ from a target network are periodically copied from $\mathit{\Theta}$ and remain constant over multiple iterations.
@@
Centralized training and decentralized execution (CTDE). In the CTDE fashion, during training the central controller can access the action-observation history of all agents and the global state, as well as the freedom to share all information between agents. While in the execution phase, each agent has its policy network to make decisions based on its individual action-observation history. Individual-global-max (IGM) is a common principle to realize effective CTDE, which asserts the consistency between joint and local greedy action selections by the joint value function $Q_{tot}$ and individual action-value function $Q_{i}(\tau_{i},u_{i})$:
@@
where $\bm{\tau}$ represents the joint action-observation history of all individual agents. Existing popular algorithms such as value decomposition networks (VDN) [2] and QMIX [3] are based on CTDE, estimating the optimal $Q_{tot}$ via combining individual action-value function $Q_{i}$. In VDN, the $Q_{tot}$ is calculated by summing the utilities of each agent as $Q_{tot}(\bm{\tau},\bm{u})=\sum_{i=1}^{n}Q_{i}(\tau_{i},u_{i})$. QMIX combines $Q_{i}$ through the state-dependent nonlinear monotonic function $f_{s}$: $Q_{tot}(\bm{\tau},\bm{u})=f_{s}(Q_{1}(\tau_{1},u_{1}),...,Q_{i}(\tau_{i},u_{i}),...,Q_{n}(\tau_{n},u_{n}))$, where $\frac{\partial f_{s}}{\partial Q_{i}}\geq 0,\forall i\in\{1,...,n\}$.
Soft decision trees (SDTs). Differentiable decision trees [41, 28] have been shown to own better expressivity compared to traditional hard decision trees. Especially, an SDT [28] has favorable transparency and performance with a binary probabilistic decision boundary at each node in reinforcement learning tasks [31, 32]. As shown in Fig. 1(a), it is different from the traditional growth of decision paths in that an SDT performs soft routing with different probabilities when given the fixed depth of the tree in advance. Given an observation $o^{t}$, each inner node $j$ calculates the probability of traversing to its left child node by
@@
Figure 1: Examples of SDT and RTC. (a) A two-level SDT. (b) Illustration of the process of a two-level RTC that receives the current individual observation $o^{t}_{i}$ and the previous hidden state $h^{t-1}_{i}$ as input at each timestamp.
@@ 2.2 Related works
Explainable reinforcement learning. In the reinforcement learning community, several prior works have investigated intrinsic methods aimed at deriving interpretable policies [29, 41, 42]. Intrinsic methods generally rely on using inherently interpretable models, such as classical decision trees and linear models, which directly represent the faithful decision-making process. The conservative Q-improvement [42] introduces an interpretable decision-tree model that learns a policy as a hard decision tree for robot navigation tasks. It can balance the performance and simplicity of the policy model only when adding new splits in a task. Differentiable decision trees methods [29, 43] allow for a gradient update over Q-learning and policy gradient algorithms by replacing the Boolean decision with a sigmoid activation function at the decision node, which improves the performance while affording interpretable tree-based policies. These methods fix their structure before the learning process rather than incrementally adding nodes, where the leaf nodes represent single features using discretization techniques to enhance interpretability. Instead of constructing a univariate decision tree, the cascading decision tree (CDT) [32], an extension based on an SDT with multivariate decision nodes, applies representation learning to decision paths to achieve richer expressive capabilities. The prior works closest to ours are SDT [28] and CDT [32], whose policies are generated by the routing probabilities at each leaf node to make decisions. Our method differs in that 1) we introduce the recurrency into SDTs for capturing the long-term condition in partially observable tasks, 2) we explicitly present the graphical decision process with a smaller depth by considering all variate at each node instead of univariate features, and 3) we interpret the role each agent plays in a team by visualizing the linear credit assignment.
@@
Cooperative multi-agent reinforcement learning. Dating back to the early stage of MARL, independent Q-learning [44] is a common method, in which each agent optimizes its policy that is independent of others. In this approach, agents observe only local information, execute their actions, and receive rewards individually, effectively treating other agents as part of the environment. However, this method lacks the theoretical convergence guarantees of standard Q-Learning due to the non-stationarity introduced by changes in other agents’ policies. By contrast, CTDE [45, 46] is an advanced paradigm for cooperative MARL tasks, which allows each agent to learn a joint action-value function via credit assignment mechanisms. VDN [2] factorizes the joint action-value function into a linear summation over individual agents. To address the VDN’s limitation of ignoring additional global state information, QMIX [3] utilizes a non-linear mixing network augmented with a state-dependent bias to estimate the joint action-value function more effectively. QTRAN [38] further relaxes the constraints on the greedy action selections between the joint and individual functions using two soft regularizations. Different from QTRAN which loses the exact IGM consistency, QPLEX [35] uses a dueling network architecture to represent both joint and individual action-value functions to guarantee the IGM property.
Our method distinguishes itself from these representative value decomposition methods in several key aspects: 1) Our method builds upon soft decision trees, which not only provide intrinsic interpretability for visualization but also preserve historical information through a recurrent structure. 2) Our mixing trees module linearly factorizes the joint action value into individual action values, achieving high scalability by fully implementing the IGM principle. This interpretable model enables us to explain not only the explicit behavior of each agent but also the behaviors of different roles within the team. 3) Our proposed method requires only lightweight linear reasoning while achieving competitive performance across a series of collaborative tasks.